OpenStudy (anonymous):

In an experiment to determine the value of pi, a cylinder is measured to have an average value of 4.25 cm for its diameter and an average value of 13.39 cm for its circumference. What is the experimental value of pi to the correct number of significant figures?

OpenStudy (asnaseer):

circumference \(C=\pi D\) where D is the diameter, so:\[\pi=\frac{C}{D}\]

OpenStudy (asnaseer):

just plug in your values for C and D and round to 2 decimal places.

OpenStudy (anonymous):

so is my answer, pi=C/D or 4.25/13.39=0.317 ? I thought it would be c muliplied by D =3.15 ?

OpenStudy (asnaseer):

in your case circumference C=13.39, and diameter D=4.25, so pi = C/D = 13.39/4.25 = 3.1506... = 3.15 (to 2 decimal places)

OpenStudy (anonymous):

Awesome. thank you

OpenStudy (asnaseer):

yw